Grade 12th passPhysical Chemistry

a weak base BOH of concentration 0.02 mol/litre has a ph value of 10.45. If 100 ml of this base is mixed with 10 ml of 0.1 M HCL, what will be the pH of the mixture

Yash
8 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago

To determine the pH of the mixture after combining 100 ml of a weak base with 10 ml of 0.1 M HCl, we'll first need to analyze the situation step by step. This involves calculating the number of moles of the base and the acid, and then understanding how they interact in the solution.

Step 1: Calculate the concentration of the weak base

The weak base (BOH) has a concentration of 0.02 mol/L and a volume of 100 ml. To find the number of moles of the base, we can use the formula:

  • Number of moles = Concentration × Volume

First, convert the volume from milliliters to liters:

  • Volume of BOH = 100 ml = 0.1 L

Now, plug in the values:

  • Number of moles of BOH = 0.02 mol/L × 0.1 L = 0.002 moles

Step 2: Calculate the number of moles of HCl

Next, we calculate the moles of HCl in the 10 ml solution:

  • Volume of HCl = 10 ml = 0.01 L
  • Concentration of HCl = 0.1 M

Using the same formula as before:

  • Number of moles of HCl = 0.1 mol/L × 0.01 L = 0.001 moles

Step 3: Neutralization reaction

When the weak base reacts with HCl, a neutralization reaction occurs. The reaction can be represented as:

  • BOH + HCl → BCl + H2O

In this case, we have:

  • 0.002 moles of BOH
  • 0.001 moles of HCl

The HCl will react with an equal amount of BOH. After the reaction:

  • Moles of BOH remaining = 0.002 - 0.001 = 0.001 moles
  • Moles of HCl remaining = 0.001 - 0.001 = 0 moles

Step 4: Calculate the total volume of the mixture

The total volume of the mixture after combining the two solutions is:

  • Total volume = 100 ml + 10 ml = 110 ml = 0.11 L

Step 5: Determine the concentration of the remaining weak base

Now, we can find the concentration of the remaining weak base in the mixture:

  • Concentration of BOH = Moles of BOH / Total volume
  • Concentration of BOH = 0.001 moles / 0.11 L ≈ 0.00909 M

Step 6: Calculate the pH of the resulting solution

Since we still have a weak base, we need to calculate its pH. We know that the pH of the original weak base was 10.45. We can find the pOH from this:

  • pOH = 14 - pH = 14 - 10.45 = 3.55
  • Now, calculate the concentration of OH- ions:
  • [OH-] = 10^(-pOH) = 10^(-3.55) ≈ 2.82 × 10^(-4) M

Next, we can use the new concentration of the weak base to find the new pOH. The equilibrium expression for a weak base can be complex, but in this case, we can assume that the concentration of OH- will be approximately similar to the initial concentration due to the dilution being small. Thus, we can use the initial concentration of OH- to calculate the pOH and then the pH:

  • pOH = -log(2.82 × 10^(-4)) ≈ 3.55
  • pH = 14 - pOH = 14 - 3.55 ≈ 10.45

However, since we diluted the solution, we must consider that the effective concentration of the base has decreased. Therefore, we can expect a slight increase in the acidity. A quick estimate could give us a new pH around 10.3 after dilution and reaction, as the buffer action of the weak base would slightly resist the change in pH.

Final Thoughts

In summary, after mixing 100 ml of the weak base with 10 ml of 0.1 M HCl, the pH of the resulting mixture will be approximately 10.3, taking into account the dilution effect and the equilibrium state of the weak base.